OpenStudy (anonymous):

Help please?(: Is this equation a trigonometric identity? sin^2θ csc^2θ=sin^2θ+cos^2θ

OpenStudy (campbell_st):

well you need a few trig identities on the right hand side you have \[\sin^2(\theta) + \cos^2(\theta) = 1\] so substitute 1 for sin^2 + cos^2 son csc^ 2 = 1 sin^2 so on the left you have \[\sin^2(\theta) \times \frac{1}{\sin^2(\theta)}\] so it should be easy to prove from here 1 = 1

OpenStudy (anonymous):

Right, and since the left side is equal to the right side, it is indeed a trig identity? @campbell_st

OpenStudy (campbell_st):

yes.... by using the appropriate substitution, you're showing 1 = 1... so the trig identity is true

OpenStudy (anonymous):

Thanks!